A saturated solution of manganese(II) hydroxide was prepared, and an acid–base titration was performed to determine its Ksp at 25 °C. The endpoint was reached when 60.00 mL of the manganese(II) hydroxide solution was titrated with 4.33 mL of 0.0030 M  HCl solution. What is the Ksp of manganese(II) hydroxide?
Answer #1

Mn(OH)2 + 2 HCl --> MnCl2 + H2O

it requires 2 moles of HCl for each mole Mn(OH)2

(1 mole Mn(OH)2 / 2 mole HCl) X (0.0030 mole HCl / 1000ml) X (4.33 ml) = 0.0065 mol

this is dissolved in (60 mL+4.33 mL = 64.33 mL)

Therefore conc of Mn(OH)2 = 0.0065/0.064 = 0.101 M

Ksp = [Mn2+][OH-]^2 = 4 * X^3 = 4 * (0.101)^3 = 0.004121 = 4.121*10^-3